Using X=x-2, Y=y+1, reduce the equation
4(x-2)^2.dy/dx=(x+y-1)^2 to the homengenous form of 1st order equation
1
Expert's answer
2020-10-19T16:34:13-0400

A function f(x,y) is said to be homogeneous of degree n if the equation "f(zx,zy)=z^nf(x,y)" .

A first‐order differential equation "M(x,y)dx+N(x,y)dy=0" is said to be homogeneous if "M(x,y)" and "N(x,y)" are both homogeneous functions of the same degree.


"4(x-2)^2\frac{dy}{dx}=(x+y-1)^2" ;

if "X=x-2" , "Y=y+1", "\frac{dY}{dX}=\frac{d(y+1)}{d(x-2)}=\frac{dy}{dx}" than:

"4(X)^2\frac{dY}{dX}=((x-2)+(y+1))^2" ;

"4X^2\frac{dY}{dX}=(X+Y)^2" ;

"(X+Y)^2dX-4X^2dY=0" .

where

"M(X,Y)=(X+Y)^2" and "N(X,Y)=-4X^2" are homogeneous functions of the same degree (namely, 2), because:

"M(zX,zY)=(zX+zY)^2=z^2(X+Y)^2" and

"N(zX,zY)=-4(zX)^2=z^2(-4X^2)" .

Answer: the reduced equation is "(X+Y)^2dX-4X^2dY=0" .